# Challenge LLMs to Reason About Reasoning: A Benchmark to Unveil Cognitive Depth in LLMs

### 28 Dec 2023 v1

@@ 5 Discussion and Limitations
Reflecting on Goodhart’s law, which states that ’When a measure becomes a target, it ceases to be a good measure,’ it’s evident that the ’reason about reasoning’ paradigm is not immune to this phenomenon. This paradigm, like any other, can be targeted for optimization, as illustrated in Section-4.3. This is particularly pertinent for static benchmarks where in-domain augmentation and overfitting are feasible. Nonetheless, we contend that our evaluation paradigm presents a greater challenge to overfitting compared to others, owing to its demand for a comprehensive understanding of the problem within a broad error space. While the current DiagGSM8K benchmark is constructed from incorrect solutions generated by the MetaMath-7B model, its scope for expansion is vast. Future versions could incorporate solutions from humans, various models, and even across different languages, enriching its complexity and utility.
Utilizing DiagGSM8K for Model Evaluation Without Manual Labelling
The complexity of the diagnostic process in DiagGSM8K necessitated manual labelling of error reasons in this study to ensure accuracy. However, it is conceivable to use GPT4 for scoring error reasons, with the evaluation of solution correctness and first error step being automated. Table-LABEL:tab:confusionMatrix displays a confusion matrix based on GPT4’s labelling of 100 sampled evaluation instances. While GPT4 exhibits a high true-positive rate, it also has a notable false-positive rate. Its overall accuracy in diagnosing incorrect error reasons aligns closely with its performance in providing error reasons, as shown in Figure-4. Consequently, we advise caution when using GPT4 for automated evaluation. Furthermore, we encourage large tech companies, which have the resources to bear manual labelling costs, to release open-source manual labelling results when publishing findings using DiagGSM8k.

### 20 Jan 2024 v2

@@ 5 Ethical Discussion and Limitations
Reflecting on Goodhart’s law, which states that ’When a measure becomes a target, it ceases to be a good measure,’ it’s evident that the ’reason about reasoning’ paradigm is not immune to this phenomenon. This paradigm, like any other, can be targeted for optimization, as illustrated in Section-4.3. This is particularly pertinent for static benchmarks where in-domain augmentation and overfitting are feasible. Nonetheless, we contend that our evaluation paradigm presents a greater challenge to overfitting compared to others, owing to its demand for a comprehensive understanding of the problem within a broad error space. While the current MR-GSM8k benchmark is constructed from incorrect solutions generated by the MetaMath-7B model, its scope for expansion is vast. Future versions could incorporate solutions from humans, various models, and even across different languages, enriching its complexity and utility.
Does MR-GSM8k Mandates Human Labelling?
For the sake of rigorousness, every error reason from evaluated models with correct first error step predicted is examined manually in this work. However, this does not imply that MR-GSM8K necessitates a manual labelling for every evaluation. We would like to emphasize that MR-Score is consist of three sub-metrics and error reason is (only) one of the evaluation criteria. Similar to the translation tasks where one expression in one language might corresponds to many variations in another language, it is likewise difficult to come up with an automatic evaluator that scores the error reason perfectly. Nonetheless, this would not undermine the arguments we contend, nor would it affect the cognitive deficiencies unveiled by this metrics. To the best of our knowledge, GPT4 has been the most popular choice for being an automatic evaluator across different metrics (Zheng et al., 2023; Liu et al., 2023). In Appendix-B we empirically demonstrated that GPT4 is able to serve as a decent automatic evaluator that the final MR-Score calculated based on its labelling results is close to that of the manual labelling results.
6 Conclusion
Throughout this paper, we have delved into the inadequacies of prevalent math reasoning benchmarks and introduced a pioneering evaluation paradigm that compels models to engage in meta-reasoning. Our empirical findings showcase that the benchmark, developed under this novel paradigm, stands out in its ability to differentiate between models and uncover their various deficiencies. This has been particularly evident in the struggles of state-of-the-art language models, which, when confronted with our benchmark, have exposed significant shortcomings inherent in the current training methodologies. These revelations advocate a critical reevaluation of existing training and evaluation practices in the realm of large language models.
In advocating for the widespread adoption of our ’reason about reasoning’ evaluation paradigm, we urge researchers to adapt and broaden other reasoning benchmarks in a similar vein. Such transformation is vital not only for a more rigorous assessment of LLMs but also for fostering a deeper and more holistic understanding of these models’ capabilities. By pushing the boundaries of conventional benchmarks and embracing this innovative approach, we can pave the way for more insightful, nuanced evaluations that truly challenge and advance the frontiers of AI research.
Appendix A Annotation Details for MR-GSM8k
In this section we will give a more thorough description of the dataset construction process. We will first present the annotation procedure adopted by our annotators and then describe the challenges faced during such process.
A.1 Annotation Procedure
Given a question and its ground truth solution, the annotator is supposed to first read the question and ground truth solution carefully and make sure he/she comprehend the question thoroughly. Then the annotator is tasked to label the following fields sequentially:
First Error Step: This attribute is applicable for solutions with either an unmatched final output or a matched final output underpinned by flawed reasoning. Annotators identify the initial step in the reasoning process where the logic deviates from correctness. In line with the approach of Lightman et al., 2023, we dissected GSM8K solutions into discrete steps, each marked by a newline character, and indexed them accordingly. Each step is then categorized as positive, neutral, or negative. Positive and neutral steps represent stages in the reasoning process where the correct final output remains attainable, whereas negative steps indicate a divergence from the path leading to the correct solution.
A.2 Annotation Challenges
The annotation task turns out to be more difficult than we expect at the beginning. The challenge comes from several sources: First, the language barrier hinders the non-native speaker to understand the meaning of the question. Second, the labelling task requires the annotator to read the question, the ground truth solution and the model solution before judging the correctness of the reasoning process. If the reasoning process is problematic, the annotator is further required to find out the first error step and reason about why the model made such mistake. To reason along the path of the model solution and figure out why the error occurred in the perspective of the model and verbalize all the above in the error reason is quite time-consuming and an exhausting task. Third, some questions are ambiguous in its wording and allows for multiple different interpretations. The backward reasoning transformation of the original problem exacerbated this problem.
To help lower the difficulty of labelling, we translated the problems and solutions into Chinese with the help from ChatGPT (OpenAI, 2022). We noticed that ChatGPT occasionally would made some translation errors such as missing critical information and misinterpret the original text, we therefore enclosed both the original problem-solution pairs and the translated ones for reference when in doubt (See Figure-7 for a full example). The solution correctness, first error step and error analysis are then collected from the annotated dictionary and translated back to English.
A.3 Annotation Quality Control
The annotators are selected based on their labelling performance on a balanced small hold-out problems set consist of 50 questions. For every problems in the Mr-GSM8K, they have been through several examinations before used in the evaluation process: First, every question is labelled twice by different annotators. Inconsistent questions will be singled out to be labelled by the quality control labeller. Besides the double annotation, 50 percent of the annotated problems are sampled out for quality control in the second round of verification. During the evaluation process, all the questions that has a matching correctness prediction and error steps are manually examined by the authors of this work for its error reasons. Questions with incorrect error steps or reasons are cleaned up in this final stage.
Figure 7: To lower the requirements and difficulties for labelling, we used ChatGPT to translate the MR-GSM8k problems into Chinese. In case for the potential translation errors (missing information and wrong translation) we included the original English text for reference when in doubt.
Appendix B Design thinking of MR-Score
The MR-Score is consist of three sub-metrics corresponds to the three sequential reasoning sub-tasks. For the first solution correctness prediction, we empirically noticed that most of the evaluated language models tend to blindly classify the given solution as correct, exemplified by the low true-negative rate in Table-2. Therefore, we chose the MCC score instead of metrics like F1 or Balanced-Accuracy due to its value range. The models that have high true-positive rate but low true negative rate will have near zero score under the MCC metric. For the second and third tasks of locating first error step and elucidating error reason, we chose the simple accuracy metric. One of the reason is that locating the first error step is a multi-class classification problem and it is difficult to have large prediction bias while at the same time scores high accuracy. Similarly, the explaining error reason task is a free-form generation task that requires substantial understanding and a simple accuracy metric is enough to categorize the model behavior.
As to the weights given to the three metrics, they are crafted by considering the task complexity and the difference between manual labelling and auto labelling results of the error reason. As discussed in Section-5, we chose the GPT4-Turbo-1106 as our proxy evaluator and Table-4 is the results of auto-labelling vs our expert manual-labelling. It is clear that the final MR-Score calculated from manual labelling VS auto labelling are very close to each other, exhibiting the potential of GPT4 to serve as a delegate evaluator for our task.
Models Step Step+Reason/M Step+Reason/A MR-Score/M MR-Score/A
GPT4 823/1573 677/1573 732/1573 0.495 0.512
Claude2 331/1573 185/1573 224/1573 0.191 0.203
llama2-70B-MR 327/1573 99/1573 139/1573 0.105 0.118
GPT3-5 179/1573 73/1573 73/1573 0.097 0.097
MetaMath-70B 22/1573 6/1573 7/1573 0.013 0.013
Mammoth-70B 4/1573 1/1573 2/1573 0.012 0.012
WizardMath-70B 6/1573 1/1573 1/1573 0.001 0.001
Table 4: Comparison of the manual labelling results and GPT4-Turbo-1106 labelling results. Step column shows the number that each evaluated models successfully located the first error steps among incorrect solutions. Step+Reason/M stands for the manual labelling results of the error reasons where its first error step is correct. Step+Reason/A corresponds to the labelling results of GPT4-Turbo-1106. llama2-70B-MR are llama2-70B model finetuned on the GSM8k training set and its meta-reasoning augmentation by GPT4.
Table-LABEL:tab:confusionMatrix displays the confusion matrix based on GPT4’s labelling of all the error reasons. Notably, GPT4 is able to achieve 82% of overall accuracy despite a substantial higher false positive rate than false negative rate. However, we still encourage large tech companies, who have the resources to bear manual labelling costs, to release open-source manual labelling results when publishing findings using MR-GSM8k for the best of rigorousness.